OpenStudy (mackenzie6302):

Write 30% as a fraction in simplest form. 7/10 1/30 3/7 3/10

OpenStudy (snowsurf):

First how do you write 30% as a fraction?

OpenStudy (skullpatrol):

% = 1/100 30% = 30 *1/100 = 30/100